# An unexpected rational blowdown

### v1

@@ 3. Braided wiring diagrams with tangencies
Consider the Scott deformation $C^{t}$ of a germ $C$ with two irreducible components, each with a simple cusp singularity, intersecting each other with multiplicity $7$. Figure 5 shows the real part of the complex curve $C^{t}$ which is the general fiber of the Scott deformation (left), and the corresponding braided wiring diagram (right). Note the intertwined strands in the neighboorhood of the two tangencies.
@@
where $\sigma_{i}$ stands, as usual, for the positive half-twist between the $i$-th and the $(i+1)$-th strands, and $\Delta_{1,4}$ is the positive half-twist on all strands $1,\dots,4$, and we used the square brackets to emphasize the contributions from the back and the front of the diagram. The reader can write a simpler expression for the (conjugacy class of) this closed braid and verify that it is consistent with the braid one gets from the Piuseux expansion.
@@
We are now ready to analyze the Scott deformation of the germs corresponding to the family of singularities (topologically) encoded by the graphs $G_{k,n}$. We ignore the marked points for the moment, since our goal is to understand the boundary braid. The diagrams in Figures 7 and 8 represent the cases k=1k=-1 and $k=0$: note that the diagrams are slightly different because the order of the strands on the right of the diagram is affected by parity. These diagrams easily generalize to all $k$ even resp. $k$ odd: increasing $k$ to $k+2$ adds two new germ components (that is, four new strands in the diagram) and two new multipoints involving the “bottom” strands of all the components. We hope that the reader is able to see how the patterns in the pictures extend to infinite families.
@@
To be able to write formulas for the diagrams, we introduce notation for certain elements that appear in our examples, see Figure 6. The braiding will be written using the standard braid generators, where $\sigma_{i}$ denotes the positive half-twist between the $i$-th and the $(i+1)$-th strands. (We avoid the word “crossing” to avoid confusion with the intersections.) The multipoint intersection of strands $i,i+1,\dots,j$ is denoted $\operatorname{I}^{i}_{j}$; we only indicate the first and the last strand since the intersection point involves consecutive strands. For an intersection of only two strands, we also write $\operatorname{I}_{i}=\operatorname{I}^{i}_{i+1}$. We write $\operatorname{X}_{j+1,k}^{i,j}$ for a configuration of double points where parallel strands $i\dots,j$ intersect parallel strands $j+1,\dots,k$ in a square grid pattern. (Since the strands are necessarily consecutive, one of the indices is redundant and could be dropped from notation, but we found that the longer notation helps keep track of strands and relations. We will write $\operatorname{X}^{i}_{i+1,k}$ resp. $\operatorname{X}^{i,j}_{j+1}$ when there is only one strand is the top or bottom collection.) The vertical tangency involving the strands $i$ and $(i+1)$ is denoted by $\operatorname{T}_{i}$; this notation only makes sense if the two strands belong to the same component of the germ. We encode the diagrams by words in these elements, written from left to right. Exponents will mean that the subword is repeated the corresponding number of times; for example, $(\operatorname{I}^{i}_{j})^{r}$ means that there are $r$ consecutive intersection points involving all strands from $i$ to $j$. With this notation, the diagram of Figure 5 is written as $\operatorname{T}_{2}\sigma_{1}^{-1}\sigma_{3}^{-1}\operatorname{T}_{2}(\operatorname{I}_{3})^{3}\operatorname{I}^{1}_{4}$.
@@
$\begin{split}\mathbb{T}_{1,2m}\operatorname{I}^{m+1+\epsilon}_{2m-1+\epsilon}\,(\operatorname{I}^{m+1}_{2m})^{m+2}\,\operatorname{I}^{1}_{2m},\end{split}$
where $\epsilon=0$ for $k$ even, $\epsilon=1$ for $k$ odd. For k=1,0k=-1,0, the diagrams are shown in Figures 7 and 8.
Figure 7. Wiring diagram for the Scott deformation for the germ of the graph $G_{-1,1}$. Here and in subsequent figures, the strands are labeled top to bottom. Strands representing the same irreducible component of the germ are shown in the same color. Diagrams for all even $n$ follow the same pattern.
Figure 8. Wiring diagram for the Scott deformation for the germ of the graph $G_{0,1}$. Diagrams for all odd $n$ follow the same pattern.
@@
This immediately follows the structure of the Scott deformation for the given germs (see Figure 3, with the careful treatment of the braiding near the tangencies. ∎
@@ 4. Boundary-preserving moves on braided wiring diagrams
Some of our moves can be seen locally as deformations of complexifications of real arrangements. For example, the moves 4.4 and 4.7 are modelled on moving a complex line by translation near the singular point. In general however, to check that the boundary braid is preserved (up to isotopy), we check that the contribution of the local portion of the braided wiring diagram to the boundary braid on the “back” and the contribution to the braid on the “front” are unchanged by the move. For some moves this is easy to see. For example, we can isotope a strand of the braided wiring diagram using braid moves, or slide a multipoint intersection under (resp. over) a strand if the strand passes over (resp. under) all of the strands involved in the multipoint (moves 4, 5(b), and 5(c) in [CS97, section 5.6]).
@@ 4.1. Commuting elements, expanding $\operatorname{X}$ and $\mathbb{T}$
A multipoint involving $n$ consecutive strands in the diagram can split in a number of ways. It can split generically into $n\choose 2$ double points, or there can be several multipoints corresponding to the intersection of some subsets of the original set of strands, with other strands intersecting at double points. The inverse move merges two or more multipoints into one. See Figure 9. Note that no braiding is introduced between the intersections when splitting a multipoint. One can see this using a model given by complexified real lines, or by directly working with braid relations for the front and back braid pushoffs. Correspondingly when merging multipoints, it is critical that there is no braiding in between the multipoints which could interfere with deforming them to a single multipoint. In our notation, when a multipoint involving strands from $i$ to $k$ splits into a multipoint of intersection of strands from $i$ to $j$ and the other of strands from $j+1$ to $k$, $i\leq j\leq k$, with the pairs of strands from these different groups meeting at double points:
(4.4) $\operatorname{I}^{i}_{k}=_{\partial}\operatorname{I}^{i}_{j}\operatorname{I}^{i}_{k}\operatorname{X}^{i,j}_{j+1,k}=_{\partial}\operatorname{X}^{i,j}_{j+1,k}\operatorname{I}^{i}_{i+k-j-1}\operatorname{I}^{i+k-j}_{k}.$
@@
for splitting strand $j$ off an intersection multipoint of strands $i$ to $k$, $i<j<k$. There are other variants for splitting into various partitions of strands.
@@
where $B=(\sigma_{n}^{-1}\sigma_{n-1}^{-1}\dots\sigma_{1}^{-1})\dots(\sigma_{2n+1}^{-1}\dots\sigma_{n}^{-1})$ is the braid where the parallel strands $1,\dots,n$ cross under the parallel strands $(n+1),\dots,2n$, and the relation above follows from the identities such as $\Delta_{1,n}B=B\Delta_{n+1,2n}$ (one can slide any braid on the strands $1,\dots,n$ past $B$ by an isotopy to get the same braid on the strands $n+1,\dots,2n$ on the other side).
@@
 $\begin{split}\text{ front }&:\Delta_{n+1,2n}^{-1}\Delta_{1,n}^{-1}\Delta_{1,n}=\Delta_{1,n}\Delta_{1,n}\Delta_{1,n}^{-1}\Delta_{n+1,2n}^{-1}\Delta_{1,n}^{-1}\\ \text{ back }&:B\Delta_{1,n}^{-1}=\Delta_{1,n}^{-1}\Delta_{1,n}B\Delta_{1,n}^{-1}.\end{split}$
@@ 5. $\mathbb{Q}$HD fillings via diagrammatic moves
Starting with the Scott deformation of the germ corresponding to the family of singularities in Theorem 1.1, we would like to obtain an arrangement with the same boundary braid, such that the number of intersection points equals the number of curves. The combinatorial type of the arrangement we seek is easy to guess: the curve $\Gamma_{0}$ of weight $7+k$ has its double point disjoint from all the other curves, and for each of the remaining curves $\Gamma_{1},\dots,\Gamma_{k+6}$, all the other curves go through the double point of $\Gamma_{i}$ with multiplicity $1$. To generate the immersed disk arrangement with the correct boundary, we use the diagrammatic moves from the previous section. The output of the moves, at the end of the procedure, will be the wiring diagrams of the desired arrangements; these are shown in Figure 16 for $k$ even and Figure 17 for $k$ odd. (The marked points for these DJVS arrangements are the intersection points in the diagram.)
@@
The proof consists of presenting a sequence of moves changing the diagrams as required; we give this sequence for $k$ even. The moves for $k$ odd are similar. For a reader willing to accept a proof by picture, the entire proof is in Figure 15 illustrating the moves connecting Figure 7 to Figure 16 for k=1k=-1; the pattern extends in a straightforward way to all even $k$. For a more rigorous argument, we spell out the required moves in formulas below.
@@
which equals $\mathbb{T}\operatorname{I}_{m+1,2m-1}\operatorname{I}_{1,m}(\operatorname{I}_{m+1,2m})^{m-4}\operatorname{X}^{2,m}_{m+1,2m}\operatorname{I}_{1,m+1}$ because the corresponding multipoints commute. Then slide the (m4)(m-4) bottom multipoints $(\operatorname{I}^{m+1}_{2m})^{m-4}$ up and combine them, one by one, with the double points on strands $2,\dots,m-3$ going down: the two multipoints in Figure 15(a) are combined with the purple and blue strands of double points, respectively. More formally, expanding $\operatorname{X}^{2,m}_{m+1,2m}=\operatorname{X}^{3,m}_{m+1,2m}\operatorname{X}^{2}_{3,m+2}$, then moving a multipoint past parallel strands by (4.7), and merging it with double points by (4.5) gives
@@
Figure 16. The pattern for arrangements producing a $\mathbb{Q}$HD filling for $k$ even (k=1k=-1 is shown).
For $k$ odd, a DJVS arrangement with the desired properties in shown in Figure 17. We omit the moves for this case, the final factorisation is as follows:
@@
As for $k$ even, one of the moves slides an intersection point past a tangency nest, adding some braiding at the beginning of the diagram. To simplify the picture, we gave an equivalent diagram with this braiding removed.
Figure 17. The pattern for arrangements producing a $\mathbb{Q}$HD filling for $n$ odd ($k=0$ is shown).
@@ 6. Constructing exotic 4-manifolds with new rational blowdowns
Consider a twice punctured torus with two meridians $\alpha_{1}$ and $\alpha_{2}$, a longitude $\beta$, and the two boundary parallel curves $\delta_{1}$ and $\delta_{2}$, as in Figure 18. For brevity, we will use the same notation for the curve and for the positive Dehn twist around this curve. By [KO08, Section 3.2], we have a relation
@@
Figure 19. Left: A configuration of embedded spheres in $E(3)$. Right: after the blow-ups, there is a subgraph $G_{0,1}$. Unlabeled vertices are understood to have framing 2-2.
@@ 7. $\mathbb{Q}$HD fillings for the two-parameter family
Then use merge move (4.5) to combine the $\operatorname{I}^{s+2}_{2}$ multipoint with crossings of the (blue) strand $(s+1)$ passing above:
@@
Slide $\operatorname{I}^{n+1}_{s}$ over intersections with red strands; it becomes $\operatorname{I}^{1}_{k+6}$. Use move (4.9) to switch $\operatorname{I}^{s+1-n}_{s+n}$ to the left of the tangency nest on the red strands:

### v2

@@ 3. Braided wiring diagrams with tangencies
Consider the Scott deformation $C^{t}$ of a germ $C$ with two irreducible components, each with a simple cusp singularity, intersecting each other with multiplicity $7$. Figure 5 shows the real part of the complex curve $C^{t}$ which is the general fiber of the Scott deformation (left), and the corresponding braided wiring diagram (right). Note the intertwined strands in the neighborhood of the two tangencies.
@@
where $\sigma_{i}$ stands, as usual, for the positive half-twist between the $i$-th and the $(i+1)$-th strands, and $\Delta_{1,4}$ is the positive half-twist on all strands $1,\dots,4$, and we used the square brackets to emphasize the contributions from the back and the front of the diagram. The reader can write a simpler expression for the (conjugacy class of) this closed braid and verify that it is consistent with the braid one gets from the Puiseux expansion.
@@
We are now ready to analyze the Scott deformation of the germs corresponding to the family of singularities (topologically) encoded by the graphs $G_{k,n}$. We ignore the marked points for the moment, since our goal is to understand the boundary braid. The diagrams in Figures 7 and 8 represent the cases k=1k=-1 and $k=0$: note that the diagrams are slightly different because the order of the strands on the right of the diagram is affected by parity. These diagrams easily generalize to all $k$ odd resp. $k$ even: increasing $k$ to $k+2$ adds two new germ components (that is, four new strands in the diagram) and two new multipoints involving the “bottom” strands of all the components. We hope that the reader is able to see how the patterns in the pictures extend to infinite families.
@@
To be able to write formulas for the diagrams, we introduce notation for certain elements that appear in our examples, see Figure 6. The braiding will be written using the standard braid generators, where $\sigma_{i}$ denotes the positive half-twist between the $i$-th and the $(i+1)$-th strands. (We do not use the word “crossing” to avoid confusion with the intersections.) The multipoint intersection of strands $i,i+1,\dots,j$ is denoted $\operatorname{I}^{i}_{j}$; we only indicate the first and the last strand since the intersection point involves consecutive strands. For an intersection of only two strands, we also write $\operatorname{I}_{i}=\operatorname{I}^{i}_{i+1}$. We write $\operatorname{X}_{j+1,k}^{i,j}$ for a configuration of double points where parallel strands $i\dots,j$ intersect parallel strands $j+1,\dots,k$ in a square grid pattern. (Since the strands are necessarily consecutive, one of the indices is redundant and could be dropped from notation, but we found that the longer notation helps keep track of strands and relations. We will write $\operatorname{X}^{i}_{i+1,k}$ resp. $\operatorname{X}^{i,j}_{j+1}$ when there is only one strand is the top or bottom collection.) The vertical tangency involving the strands $i$ and $(i+1)$ is denoted by $\operatorname{T}_{i}$; this notation only makes sense if the two strands belong to the same component of the germ. We encode the diagrams by words in these elements, written from left to right. Exponents will mean that the subword is repeated the corresponding number of times; for example, $(\operatorname{I}^{i}_{j})^{r}$ means that there are $r$ consecutive intersection points involving all strands from $i$ to $j$. With this notation, the diagram of Figure 5 is written as $\operatorname{T}_{2}\sigma_{1}^{-1}\sigma_{3}^{-1}\operatorname{T}_{2}(\operatorname{I}_{3})^{3}\operatorname{I}^{1}_{4}$.
@@
$\begin{split}\mathbb{T}_{1,2m}\operatorname{I}^{m+1+\epsilon}_{2m-1+\epsilon}\,(\operatorname{I}^{m+1}_{2m})^{m-4}\,\operatorname{I}^{1}_{2m},\end{split}$
where $\epsilon=0$ for $k$ odd, $\epsilon=1$ for $k$ even. For k=1,0k=-1,0, the diagrams are shown in Figures 7 and 8.
Figure 7. Wiring diagram for the Scott deformation for the germ of the graph $G_{-1,1}$. Here and in subsequent figures, the strands are labeled top to bottom. Strands representing the same irreducible component of the germ are shown in the same color. Diagrams for all odd $k$ follow the same pattern.
Figure 8. Wiring diagram for the Scott deformation for the germ of the graph $G_{0,1}$. Diagrams for all even $k$ follow the same pattern.
@@
This immediately follows the structure of the Scott deformation for the given germs (see Figure 3), with the careful treatment of the braiding near the tangencies. ∎
@@ 4.1. Commuting elements, expanding $\operatorname{X}$ and $\mathbb{T}$
A multipoint involving $n$ consecutive strands in the diagram can split in a number of ways. It can split generically into $n\choose 2$ double points, or there can be several multipoints corresponding to the intersection of some subsets of the original set of strands, with other strands intersecting at double points. The inverse move merges two or more multipoints into one. See Figure 9. Note that no braiding is introduced between the intersections when splitting a multipoint. One can see this using a model given by complexified real lines, or by directly working with braid relations for the front and back braid pushoffs. Correspondingly, when merging multipoints, it is critical that there is no braiding in between the multipoints which could interfere with deforming them to a single multipoint. In our notation, when a multipoint involving strands from $i$ to $k$ splits into a multipoint of intersection of strands from $i$ to $j$ and the other of strands from $j+1$ to $k$, $i\leq j\leq k$, with the pairs of strands from these different groups meeting at double points:
(4.4) $\operatorname{I}^{i}_{k}=_{\partial}\operatorname{I}^{i}_{j}\operatorname{I}^{j+1}_{k}\operatorname{X}^{i,j}_{j+1,k}=_{\partial}\operatorname{X}^{i,j}_{j+1,k}\operatorname{I}^{i}_{i+k-j-1}\operatorname{I}^{i+k-j}_{k}.$
@@
for splitting strand $j$ off an intersection multipoint of strands $i$ to $k$, $i<j<k$. There are other variants for splitting into various partitions of the set of strands.
@@
where $B=(\sigma_{n}^{-1}\sigma_{n-1}^{-1}\dots\sigma_{1}^{-1})\dots(\sigma_{2n+1}^{-1}\dots\sigma_{n}^{-1})$ is the braid where the parallel strands $1,\dots,n$ cross under the parallel strands $(n+1),\dots,2n$, and the relation above follows from the identities such as $\Delta_{1,n}B=B\Delta_{n+1,2n}$ (one can slide any braid on the strands $1,\dots,n$ past $B$ by an isotopy to get the same braid on the strands $n+1,\dots,2n$ on the other side). See Figure 14.
@@
 $\begin{split}\text{ front: }&\Delta_{n+1,2n}^{-1}\Delta_{1,n}^{-1}\Delta_{1,n}=\Delta_{1,n}\Delta_{1,n}\Delta_{1,n}^{-1}\Delta_{n+1,2n}^{-1}\Delta_{1,n}^{-1}\\ \text{ back: }&B\Delta_{1,n}^{-1}=\Delta_{1,n}^{-1}\Delta_{1,n}B\Delta_{1,n}^{-1}.\end{split}$
@@ 5. $\mathbb{Q}$HD fillings via diagrammatic moves
In this section, we give a proof of Theorem 1.1 for graphs $G_{k,1}$, postponing the general case until Section 7. Starting with the Scott deformation of the germ corresponding to the family of singularities in Theorem 1.1, we would like to obtain an arrangement with the same boundary braid, such that the number of intersection points equals the number of curves. The latter requirement holds for the combinatorial arrangement where the curve $\Gamma_{0}$ of weight $8+k$ has its double point disjoint from all the other curves, and for each of the remaining curves $\Gamma_{1},\dots,\Gamma_{k+6}$, all the other curves go through the double point of every $\Gamma_{i}$ with multiplicity $1$ for $i=1,\dots,k+6$. To generate the immersed disk arrangement with prescribed combinatorial data and the correct boundary, we use the diagrammatic moves from the previous section. The output of the moves, at the end of the procedure, will be the wiring diagrams of the desired arrangements; these are shown in Figure 16 for $k$ odd and Figure 17 for $k$ even. (The marked points for these DJVS arrangements are the intersection points in the diagram.)
@@
The proof consists of presenting a sequence of moves changing the diagrams as required; we give this sequence for $k$ odd. The moves for $k$ even are similar. For a reader willing to accept a proof by picture, the entire proof is in Figure 15 illustrating the moves connecting Figure 7 to Figure 16 for k=1k=-1; the pattern extends in a straightforward way to all odd $k$. For a more rigorous argument, we spell out the required moves in formulas below.
@@
which equals $\mathbb{T}\operatorname{I}^{m+1}_{2m-1}\operatorname{I}^{1}_{m}(\operatorname{I}^{m+1}_{2m})^{m-4}\operatorname{X}^{2,m}_{m+1,2m}\operatorname{I}^{1}_{m+1}$ because the corresponding multipoints commute. Then slide the (m4)(m-4) bottom multipoints $(\operatorname{I}^{m+1}_{2m})^{m-4}$ up and combine them, one by one, with the double points on strands $2,\dots,m-3$ going down: the two multipoints in Figure 15(a) are combined with the purple and blue strands of double points, respectively. More formally, expanding $\operatorname{X}^{2,m}_{m+1,2m}=\operatorname{X}^{3,m}_{m+1,2m}\operatorname{X}^{2}_{3,m+2}$, then moving a multipoint past parallel strands by (4.7), and merging it with double points by (4.5) gives
@@
Figure 16. The pattern for arrangements producing a $\mathbb{Q}$HD filling for $k$ odd (k=1k=-1 is shown).
For $k$ even, a DJVS arrangement with the desired properties in shown in Figure 17. We omit the moves for this case, the final factorization is as follows:
@@
As for $k$ odd, one of the moves slides an intersection point past a tangency nest, adding some braiding at the beginning of the diagram. To simplify the picture, we gave an equivalent diagram with this braiding removed.
Figure 17. The pattern for arrangements producing a $\mathbb{Q}$HD filling for $k$ even ($k=0$ is shown).
@@ 6. Constructing exotic 4-manifolds with new rational blowdowns
Consider a twice-punctured torus with two meridians $\alpha_{1}$ and $\alpha_{2}$, a longitude $\beta$, and the two boundary parallel curves $\delta_{1}$ and $\delta_{2}$, as in Figure 18. For brevity, we will use the same notation for the curve and for the positive Dehn twist around this curve. By [KO08, Section 3.2], we have a relation
@@
Figure 19. Left: A configuration of embedded spheres in $E(3)$. Right: after the blowups, there is a subgraph $G_{0,1}$. Unlabeled vertices are understood to have framing 2-2.
@@ 7. $\mathbb{Q}$HD fillings for the two-parameter family
Then use merge move (4.5) to combine the $\operatorname{I}^{s+2}_{2s}$ multipoint with crossings of the (blue) strand $(s+1)$ passing above:
@@
Slide $\operatorname{I}^{n+1}_{s}$ over the crossings with red strands; it becomes $\operatorname{I}^{1}_{k+6}$. Use move (4.9) to switch $\operatorname{I}^{s+1-n}_{s+n}$ to the left of the tangency nest on the red strands: